OpenStudy (anonymous):

Integration help.. how does (d^4)y/dx^4 = w / EI gives ( when integrated 4 times) y(x)=c1 + c2x +c3x^2 + c4x^4 + (w/IE)x^4 I just can't see it... any help>?

OpenStudy (fibonaccichick666):

can you define w, E, l ?

OpenStudy (fibonaccichick666):

are they constants?

OpenStudy (anonymous):

I can't its for beam deflection and this is just a general equation, but the question is based on it. w is a constant, so is e and I , just not defined in my case

OpenStudy (fibonaccichick666):

kk, constant is all I need to know

OpenStudy (fibonaccichick666):

so, if you have \[\int x\] what does that equal?

OpenStudy (fibonaccichick666):

xdx sorry

OpenStudy (fibonaccichick666):

@MarcLeclair

OpenStudy (anonymous):

its x^2/2

OpenStudy (anonymous):

+ c for undefined integrals

OpenStudy (fibonaccichick666):

ah, but you are missi-

OpenStudy (fibonaccichick666):

yes, plus c. good

OpenStudy (fibonaccichick666):

now how about \[\int 1 dx\]

OpenStudy (anonymous):

mhmmm x + c

OpenStudy (anonymous):

sorry taking long to answer, doing 2 things at the sametime

OpenStudy (fibonaccichick666):

good, now \[\int (x+c)dx\]

OpenStudy (fibonaccichick666):

stay with me for 5 minutes and you'll be done

OpenStudy (anonymous):

x+c will give you x^2/2 + c and x + c

OpenStudy (fibonaccichick666):

well not quite

OpenStudy (fibonaccichick666):

so x^2/2 yes, but just x, no

OpenStudy (fibonaccichick666):

you need to integrate \[\int C~dx\]

OpenStudy (fibonaccichick666):

what happens to a constant when we integrate?

OpenStudy (anonymous):

c would be a constant, like 1 , so x + c no?

OpenStudy (fibonaccichick666):

no, think, if you have \[\int 3x ~dx\] What do you get?

OpenStudy (anonymous):

i would get 3x^2/2

OpenStudy (fibonaccichick666):

good, so what did you do with the 3?

OpenStudy (anonymous):

just factored it out

OpenStudy (fibonaccichick666):

then multiplied it back right?

OpenStudy (anonymous):

yeah haha

OpenStudy (fibonaccichick666):

so you can't just drop your constants when you integrate something like C

OpenStudy (fibonaccichick666):

so, what is the integral of C dx now?

OpenStudy (anonymous):

it would be cx

OpenStudy (fibonaccichick666):

cx + what?

OpenStudy (anonymous):

cx + c^2 righjt?

OpenStudy (anonymous):

c ( x + c)

OpenStudy (fibonaccichick666):

no, it's Cx +(some new constant)

OpenStudy (fibonaccichick666):

we add a different constant each time

OpenStudy (anonymous):

wait, because we have c(int x dx ) wouldnt it be c ( x^2/2 + c)

OpenStudy (fibonaccichick666):

nah, we did \[C\int 1dx\]

OpenStudy (fibonaccichick666):

and remember we do a new constant every time, also you have to distribute to every value inside the parentheses

OpenStudy (fibonaccichick666):

That's why we end up with \(C_1,C_2,C_3\)

OpenStudy (anonymous):

yeah so I first integrate 1, it should give me x + c , however in this case its always multiplied ( ie x*c2 ) which is confusing to me

OpenStudy (fibonaccichick666):

noooo you do not get x+c

OpenStudy (fibonaccichick666):

you get \(C_1x+C_2\) or \(Cx+a\) whatever you wanna add, but you have to keep your constant

OpenStudy (anonymous):

mhm so if I start with int ( d^4y/ dx^4) i would then get c1 d^3y/d^3x then its going to be c1x + c2? and so on?

OpenStudy (fibonaccichick666):

just like when you had 3, you still got 3/2x^2 +c the three didn't go away

OpenStudy (fibonaccichick666):

uhh, you don't go about it that way

OpenStudy (fibonaccichick666):

but I thik you have the right idea, but your first constant is your w/el thingy

OpenStudy (fibonaccichick666):

So its just a constant call it Z yea? you do \[\int Zdx=ZX+C_1\]

OpenStudy (anonymous):

\[𝑦(𝑥) = 𝑐1 + 𝑐2𝑥 + 𝑐3𝑥 2 + 𝑐4𝑥 3 + 𝑤0 24𝐸𝐼 𝑥 4\] that would be the complete answer.

OpenStudy (anonymous):

woahhh sorry formating. \[𝑦(𝑥) = 𝑐1 + 𝑐2𝑥 + 𝑐3𝑥 2 + 𝑐4𝑥 3 + 𝑤 /24𝐸𝐼 (𝑥 4)\]

OpenStudy (fibonaccichick666):

close

OpenStudy (fibonaccichick666):

h/o

OpenStudy (fibonaccichick666):

yes, that is correct

OpenStudy (fibonaccichick666):

(I was looking at the original) Now they just decided I don't want there to be a 24 near my big fraction constant so I undistribute the 24 and have \((C_4+Z)\)

OpenStudy (fibonaccichick666):

\((C_4+Z)x^4\) ***

OpenStudy (anonymous):

mhm,. the thing that throws me off is the fact that we started with the constant w/IE and we end up with x^4 in the same term. To me it goes something like: int ( d^4y/dx^4) = w/IE , I would use DE to solve that, no?

OpenStudy (fibonaccichick666):

no, it's an integral on both sides

OpenStudy (fibonaccichick666):

it's \[\frac{d^3y}{dx^3}=\int Zdx\]

OpenStudy (fibonaccichick666):

then \[\frac{d^2y}{dx^2}=\int \int Zdx\]

OpenStudy (fibonaccichick666):

should be another dx there, but you get the gist

OpenStudy (anonymous):

I see... then that would give you something like Z ( x + c) for the first integral? and so on?

OpenStudy (anonymous):

I do get the gist though :)

OpenStudy (fibonaccichick666):

distribute the Z

OpenStudy (fibonaccichick666):

That is very important to do

OpenStudy (fibonaccichick666):

in order to get that particular general form that is

OpenStudy (anonymous):

so then Im left with zx + zc where zc can be C3 so then id have int zx + c3. this would give me then z(x+ c) + c3(x+c) and so on?

OpenStudy (anonymous):

x^2/2*

OpenStudy (fibonaccichick666):

you are adding too many constants

OpenStudy (fibonaccichick666):

you need to add one per integration or combine them when possible

OpenStudy (fibonaccichick666):

and no, you can't get rid of z

OpenStudy (fibonaccichick666):

so first integral yeilds \[zx+c_1\] second we get \[zx^2/2+c_1x+c_2\]

OpenStudy (fibonaccichick666):

now you do 3rd and fourth

OpenStudy (anonymous):

thirs would be zx^3/6 + c1x^2/2+c2x + c3

OpenStudy (fibonaccichick666):

good

OpenStudy (fibonaccichick666):

and fourth?

OpenStudy (anonymous):

zx^4/24 + c1x^3/6+ c2x^2/2+ c3x + c4.

OpenStudy (anonymous):

Thanks for your patience man, super appreciated.

OpenStudy (fibonaccichick666):

now, absorb those extra numbers into the respective constants and separate the x^4 constants and you are done

OpenStudy (fibonaccichick666):

np, happy you got it